Richard can build a dog house in 3 days. Art can build a doghouse in 6 days. If they work together, how long will it take to build a doghouse?
Answer by richwmiller(17219)

1/3+1/6=1/x
2/6+1/6=1/x
1/2=1/x
x=2
check
2/3+2/6=1
ok